OpenStudy (anonymous):

What is the equation of the axis of symmetry for the function shown below? y-2=-(x+5)^2 A. x = -5 B. x = 5 C. y = 2 D. x = 2

OpenStudy (anonymous):

The axis of symmetry goes through the vertex vertically.

OpenStudy (anonymous):

This equation is in vertex form. Can you tell what the vertex is?

OpenStudy (anonymous):

(5,-2)

OpenStudy (anonymous):

Close! (-5,2) actually.

OpenStudy (anonymous):

Recall that vertex form is: \[y-k = C(x-h)^2\] Where (h,k) is the vertex.

OpenStudy (anonymous):

So your version: y-2=-(x+5)^2 Is equivalent to: y - 2 = -1(x-(-5))^2 Therefore h = -5, and k = 2 So (-5,2) is the vertex.

OpenStudy (anonymous):

Now, what is the equation of a line that is vertical (up and down) that goes through (-5,2)

OpenStudy (anonymous):

y=2

OpenStudy (anonymous):

y = 2 is a horizontal line that goes through the y axis at y = 2.

OpenStudy (anonymous):

a vertical line has the form x = h where h is the point on the x axis that the line intersects.